Renata is the sole proprietor of a company with the following balance sheet:

Assets

Liabilities

Cash

$   100,000

Accounts payable

$     50,000

Inventory

300,000

Loan payable

100,000

Plant and equipment, net

600,000

Total liabilities

150,000

________

Capital

    850,000

Total assets

$1,000,000

Total liabilities and capital

$1,000,000

The cash and inventory are carried at fair value, and plant and equipment has a fair value of $650,000.

Renata enters into a partnership with Santiago. Renata contributes her company, and the partnership assumes the company’s liabilities. Santiago contributes cash of $250,000. The partners agree to share capital and profits in a 3:2 ratio.

Required

Prepare the balance sheet of the partnership at the date of formation using:

a.            The bonus approach

b.            The goodwill approach
Answer #1

Valuation of Business:

Particulars Amount
Cash 100000
Inventory 300000
Plant and equipment, net   650000
Total 1050000
Accounts Payable 50000
Loan Payable 100000
Net Assets Taken over 900000

2) Ratio of sharing capital:

Renata : 3

Santiago : 2.

Goodwill brought in by Santiago will be calculated as follows:

Required capital contribution for 2/5th share = 900000/3*2 = $ 600000

Capital brought in by Santiago = $ 250000

Goodwill brought in by Santiago = $ 350000

Balance Sheet Under Goodwill Approach:

Assets Amount Liabilities Amount
Cash 350000 Accounts Payable 50000
Inventory 300000 Loan Payable 100000
Plant and Equipment, Net 650000 Capital
Goodwill 350000 Renata 900000
Santiago 600000
Total 1650000 Total 1650000

Balance Sheet Under Bonus Approach:

In this method, goodwill does not appear in the books of account and bonus is shared between all the partners in the new ratio by all partners.

Particulars Amount
Cash 350000
Inventory 300000
Plant Equipment, Net 650000
Total 1300000
Accounts Payable 50000
Loan Payable 100000
Net Assets 1150000
Reneta (3) 690000
Santiago (2) 460000

Passing the entries for capital brought in by Santiago and allocating share in capital:

Cash A/c________________Dr. 250000

Reneta's Capital A/c_______Dr. 210000

To Santiago's Capital A/c - 460000

(Being capital brought in by incoming partner)

Assets Amount Liabilities Amount
Cash 350000 Accounts Payable 50000
Inventory 300000 Loan Payable 100000
Plant and Equipment, Net 650000 Capital
Renata 690000
Santiago 460000
Total 1300000 Total 1300000